OpenStudy (anonymous):

Differentiate Implicitly. x+xy-2x^3=2 Can someone please show me how?

OpenStudy (freckles):

@eggshell do you know how to differentiate x and y with respect to x?

OpenStudy (freckles):

\[\frac{d}{dx}x=\frac{dx}{dx}=1 \\ \frac{d}{dx}y=\frac{dy}{dx} =y' \]

OpenStudy (freckles):

basically if you know this you are golden plus all the rules like product, sum, constant multiply rule, constant rule, and chain rule, quotient rule...

OpenStudy (freckles):

\[\frac{d}{dx}(x+xy-2x^3)=\frac{d}{dx}(2)\] on the left use sum rule on the right use constant rule that is \[\frac{d}{dx}(x)+\frac{d}{dx}(xy)-\frac{d}{dx}(2x^3)=0\]

OpenStudy (freckles):

so can you tell me which of the 3 derivatives there you are having issues with?

OpenStudy (anonymous):

@freckles yes I know how to differentiate but not the xy

OpenStudy (freckles):

x*y is a product

OpenStudy (freckles):

you need to use product rule

OpenStudy (freckles):

\[\frac{d}{dx}(xy)=y \frac{d}{dx}x+x \frac{d}{dx}y\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

wait but that gives me x+y so it gives me (1+x+y-6x^2) the book says the answer is (6x^2-y-1)/x

OpenStudy (freckles):

no no...

OpenStudy (freckles):

did you use this: \[\frac{d}{dx}x=\frac{dx}{dx}=1 \\ \frac{d}{dx}y=\frac{dy}{dx} =y' \]

OpenStudy (freckles):

\[\frac{d}{dx}(xy)=y \frac{dx}{dx}+x \frac{dy}{dx}=y \cdot 1 +x \cdot y'=y+xy' \]

OpenStudy (freckles):

\[\frac{d}{dx}(x)+\frac{d}{dx}(xy)-\frac{d}{dx}(2x^3)=0 \\ 1+y+xy'-6x^2=0\] now you solve that for y'

OpenStudy (anonymous):

the d/dx and dy/dx thing confuse can you explain the difference?

OpenStudy (freckles):

d/dx means we are taking the derivative with respect to x dy/dx means we have taken the derivative of y with respect to x

OpenStudy (freckles):

like you don't write d/dx by itself like you have to have d/dx (some function) in order to take derivative of that function with respect to x

OpenStudy (freckles):

d/dx (y) =dy/dx=y'

OpenStudy (freckles):

maybe you prefer that latter notation (xy)'=(x)'y+x(y)'=1y+xy' since (x)'=1 and (y)'=y'

OpenStudy (anonymous):

No I understand it but isn't y' just one because you bring down the exponent? Or should I just leave it as y' when doing implicit differentiation?

OpenStudy (anonymous):

and what does with respect to x mean?

OpenStudy (freckles):

you put y' when you find the derivative of y with respect to x

OpenStudy (freckles):

and no y' doesn't equal 1 unles y=x

OpenStudy (freckles):

but we are not given y=x

OpenStudy (freckles):

y is a function of x

OpenStudy (anonymous):

But why does this example make it look like its the exact same thing as in derive both x and y exactly the same

OpenStudy (freckles):

they used chain rule for y part and power rule for x part

OpenStudy (freckles):

they used that derivative wrt to x of y is y'

OpenStudy (freckles):

they are treating y and x no different than i have

OpenStudy (freckles):

\[\frac{d}{dx}(4x^2)=4 \cdot 2 x^{2-1} \frac{dx}{dx}=8x(1)=8x \\ \frac{d}{dx}(-2y^2)=-2 \cdot 2y^{2-1} \frac{dy}{dx}=-4y \frac{dy}{dx}\]

OpenStudy (freckles):

and I have said here: \[\frac{d}{dx}(xy)=y \cdot \frac{dx}{dx}+x \cdot \frac{dy}{dx} \text{ by product rule } \\ \frac{d}{dx}(xy)=y \cdot 1 +x \frac{dy}{dx}=y+x \frac{dy}{dx}\]

OpenStudy (freckles):

just like when i differentiated y above i wrote y'

OpenStudy (freckles):

shouldn't be any different when differentiate y elsewhere

OpenStudy (anonymous):

ahhh I get it now ! it says to also solve the equation for y as a function of x and to find dy/dx. what is that asking em to do?

OpenStudy (freckles):

\[1+y+xy'-6x^2=0\] it wants you to solve this for y' put everything on the other side by addition or subtraction (except the xy' term) then divide both sides by x.

OpenStudy (freckles):

this will give you dy/dx as a function of x

OpenStudy (freckles):

oh and it looks like they also want you to solve your first equation for y too

OpenStudy (anonymous):

how would I do that second part?